OpenStudy (anonymous):

Please Help! Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = one fourth; y intercept at (0, 1) slope = one fourth; y intercept at (0, -1) slope = -4; y intercept at (0, 1)

OpenStudy (anonymous):

Im Okay .... whats the question .?

OpenStudy (anonymous):

I have a few. I think I know how to do them but I just want to make sure. :) heres 1. Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = one fourth; y intercept at (0, 1) slope = one fourth; y intercept at (0, -1) slope = -4; y intercept at (0, 1)

OpenStudy (anonymous):

x-intercept occurs where y = 0. y-intercept occurs where x = 0. 1. x - 4y = 4 x-intercept is (4,0) y-intercept is (0,-1) 2. x - 2y = 8 x-intercept is (8,0) y-intercept is (0,-4) I hope this helps!

OpenStudy (anonymous):

Thanks! Thats what I thought the answer was, but I wasn't sure. :)